OpenStudy (anonymous):

The function g is given by g(x)=eax+f(x) for all real numbers, where a is a constant. Find g ′(0) and g ″(0) in terms of a. Show the work that leads to your answers.

OpenStudy (solomonzelman):

I think that you left out some conditions from the questions. Let's suppose that, \(\color{#000000 }{ \displaystyle f'(0) =-5 }\) \(\color{#000000 }{ \displaystyle f''(0) =1 }\) Then, you can differentiate the function and get: \(\color{#000000 }{ \displaystyle g'(x)=ae^{ax}+f'(x) }\) \(\color{#000000 }{ \displaystyle g''(x)=a^2e^{ax}+f''(x) }\) Now, let's plug in \(\small x=0\) to solve for \(\small f'(0)\) and \(\small f''(0)\). \(\color{#000000 }{ \displaystyle g'(0)=a+f'(0) }\) \(\color{#000000 }{ \displaystyle g''(0)=a^2+f''(0) }\) Now, you can plug in the given values for \(\small f'(0)\) and \(\small f''(0)\). \(\color{#000000 }{ \displaystyle g'(0)=a-5 }\) \(\color{#000000 }{ \displaystyle g''(0)=a^2+1 }\)

OpenStudy (solomonzelman):

I thought of using greek letters, but I threw out the fenciness to obtain understanding